40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 18:26:05 2019, gautam, Update, IOO, IPPOS recommissioned IMG_7330.JPGScreenshot_from_2019-02-20_19-43-27.png
    Reply  Fri Feb 22 12:19:46 2019, gautam, Update, IOO, TT coil driver Vmon 
Message ID: 14467     Entry time: Wed Feb 20 18:26:05 2019     Reply to this: 14469
Author: gautam 
Type: Update 
Category: IOO 
Subject: IPPOS recommissioned 

I've suspected that the TTs are drifting significantly over the course of the last couple of days, because despite repeated alignment efforts, the AS beam spot has drifted off the center of the camera view. I tried looking at IPPOS, but found that there was no data. Looking at the table, the QPD was turned backwards, and the DAQ cable wasn't connected (neither at the PD end, nor at 1Y2, where instead, a cable labelled "Spare QPD" was plugged in). Fortunately, the beam was making it out of the vacuum. So as to have a quantitative diagnostic, I reconnected the QPD, turned it the right way round, and adjusted the steering onto it such that with the AS spot on the center of the CCD monitor, the beam is also centered on the QPD. The calibration is uncertain, but at least we will be able to see how much the spot drifts on the QPD over some days. Also, we only have 16 Hz readback of this stuff.

I leave it to Chub to take the high-res photo and update the wiki, which was last done in 2012.


Already, in the last ~1 hour, there has been considerable drift - see Attachment #2. The spot, which started at the center of the CCD monitor, has now nearly drifted off the top end. The ITMX and BS Oplev spots have been pretty constant over the same timescale, so it has to be the TTs?

Attachment 1: IMG_7330.JPG  2.080 MB  | Hide | Hide all
IMG_7330.JPG
Attachment 2: Screenshot_from_2019-02-20_19-43-27.png  113 kB  Uploaded Wed Feb 20 19:43:55 2019  | Hide | Hide all
Screenshot_from_2019-02-20_19-43-27.png
ELOG V3.1.3-